the tens digit of a two digit number is 5 more than the units digit. five times the units digit is 1 less than twice the tens digit. find the number
Answer by mananth(16949)

the tens digit of a two digit number is 5 more than the units digit. five times the units digit is 1 less than twice the tens digit. find the number
Let the number be xy
X=y+5
5y=2x-1
Re write these eequations
x-y =5------------1
-2x+5y = -1---------2
Multiply equation 1 by 2 and add it to equation 2
2x-2y=10
+3y =9
y=9/3
y=3
So x=8
The number is 83